OpenStudy (anonymous):

Find the area between the curves y^2=x and 3y^2=4-x

OpenStudy (raden):

by integral ?

OpenStudy (anonymous):

yes by integral..

OpenStudy (raden):

u have to determine the intervals for integral. they are can be found by intersection of that curves

OpenStudy (raden):

y^2 = x -------> x=y^2 ... (1) 3y^2=4-x ----> x=4-3y^2 ... (2) subtract (1) and (2) : 4y^2 - 4 = 0 y^2 - 1 = 0 (y+1)(y-1) = 0 for zeroes, y = -1 or y = 1

OpenStudy (raden):

so, the area(A) = int [-1,1] (4-3y^2-y^2) dy A = int [-1,1] (4 - 4y^2) dy solve this integral ....

OpenStudy (raden):

miss "A" of the last, a typo

OpenStudy (anonymous):

I solved it , the Area = 16/3 square units .. ?

OpenStudy (raden):

yes, i got it like u

OpenStudy (anonymous):

thanks man :)

OpenStudy (raden):

very welcome